# Emergence of Vorticity and Viscous Stress in Finite Scale Quantum Hydrodynamics

### v1

@@ Abstract
The Madelung equations offer a hydrodynamic description of quantum systems, from single particles to quantum fluids. In this formulation, the probability density is mapped onto the fluid density and the phase is treated as a scalar potential generating the velocity field. As examples of potential flows, quantum fluids described in this way are inherently irrotational, but quantum vortices may arise at discrete points where the phase is undefined. In this paper, starting from this irrotational description of a quantum fluid, a coarse-graining procedure is applied to arrive at a macroscopic description of the quantum fluid in which the role of velocity is played by a Favre average of the microscopic velocity field, allowing for finite vorticity at any point in the fluid. It is shown that this vorticity obeys a similar equation to the vorticity equation in classical hydrodynamics and includes a vortex-stretching term. This coarse-graining procedure also gives rise to novel stress terms in the fluid equations, which in the appropriate limit appear analogous to artificial viscous stresses from computational fluid dynamics.
@@
Fluid turbulence is known to be a ubiquitous non-equilibrium phenomenon arising in fluids when the kinetic energy injected at large scales dominates the dissipation of energy at small scales due to viscosityMcComb (1990, 1995); Sreenivasan (1999); Eyink and Sreenivasan (2006). While first identified and extensively studied in classical fluids, turbulence has also been observed in quantum fluids, including both superfluids and ultra-cold atomic gasesVinen and Niemela (2002); Barenghi and Donnelly (2009); Vinen (2010); Paoletti and Lathrop (2011); Tsubota (2014); Barenghi et al. (2014). While there is no consensus on the precise definition of turbulence, it is characterized by spatially complex and temporally aperiodic flow fields and involves processes on many length scales. In three-dimensions, it is well established that turbulence involves a cascade of energy from large to small length scalesKolmogorov (1991); Falkovich (2009); Dubrulle (2019); Carbone and Bragg (2020). While the details of this process are not fully understood, there is significant evidence that, in classical turbulence, this cascade is related to the non-linear phenomena of strain self-amplification and vortex stretchingCarbone and Bragg (2020). In quantum turbulence, this cascade has a remarkably similar behavior at large length scales, with the Kolmogorov spectrum being observed in superfluid 4He and simulations using the Non-linear Schrödinger equation (NLSE)Paoletti and Lathrop (2011). However, in contrast to the case of classical turbulence where the cascade is related to the dynamics of a continuous vorticity field, the cascade in quantum turbulence is caused by the interaction of a discrete number of quantized vortices and transverse modes (Kelvin waves) which propagate along these vorticesBarenghi and Donnelly (2009). While this similarity between classical and quantum turbulence is not unexpected based on intuition from the correspondence principlePaoletti and Lathrop (2011), a full theory establishing the rigorous connection between these two regimes is lacking, due in part to the lack of a single governing equation used to describe quantum turbulence at all length scales.
@@
In contrast to the study of classical turbulence, whose phenomenology is believed to emerge from the dynamics governed by a single equation, the Navier-Stokes equation, current descriptions of quantum turbulence require a hierarchy of modelsBarenghi et al. (2014). The microscopic behavior of the quantum fluid is modeled using the NLSE, or appropriate generalizations to capture strong interactions; the mesoscale behavior is described by vortex-filament models; and macroscopic dynamics are described using the Hall–Vinen–Bekharevich–Khalatnikov (HVBK) modelBarenghi et al. (2014); Zhang et al. (2023). The purpose of the present paper is to demonstrate that a suitable coarse-graining of the microscopic quantum fluid equations does in fact yield fluid equations very similar in form to the Navier-Stokes equation. Specifically, it is shown that the coarse-grained velocity field can possess finite vorticity, which obeys a similar equation to the vorticity equation of classical fluids, and that viscous stress terms also emerge as a consequence of this coarse-graining. The emergence of a novel viscous stress term due to integrating out small scale degrees of freedom has similarities with a recent work which demonstrated a relationship between decoherence and viscosityZhou et al. (2025).
@@
The particular coarse-graining procedure applied in this paper is based on finite scale theory, a method previously studied in the context of classical continuum dynamicsMargolin and Rider (2002); Margolin (2009, 2014, 2019). The motivation for finite scale theory in classical continua is based on the fact that in order to solve many problems in classical continuum dynamics, the continuum equations are almost always approximated by discretizing the domain and forming a mesh, represented by a finite number of points in the original space. In the context of Lagrangian hydrocodes, this discretization amounts to an averaging of the continuum variables over finite size cells. However, these averaged quantities are not necessarily solutions to the original partial differential equations governing the dynamics of the continuum. Finite scale theory provides a connection between the coarse-graining of the continuous domain and the differential equations describing the averaged quantities.
Here I apply a generalized version of finite scale analysis to the Madelung equations of quantum hydrodynamics. I find that, through this coarse-graining procedure new terms emerge at finite length scales, and, moreover, the Favre averaged velocity field describing the evolution of the quantum fluid can possess finite vorticity.
@@ I Madelung Equations
$\frac{1}{\rho}\boldsymbol{\nabla}\cdot\boldsymbol{\sigma}+\textbf{f}=\gamma\boldsymbol{\nabla}\left(\frac{\nabla^{2}\sqrt{\rho}}{\sqrt{\rho}}\right).$
@@ II Generalized Finite Scale Analysis
where $f$ is a normalized distribution describing the details of our coarse-graining procedure. The procedure employed in previous finite scale analyses would correspond to the choice of a uniform box distributionMargolin and Rider (2002); Margolin (2009); however, in principle, many other choices could yield similar results. For simplicity, we will assume that $f$ is a separable distribution, $f(\textbf{x})=f_{1}(x_{1})f_{2}(x_{2})f_{3}(x_{3})$ and that each function $f_{i}(x)$ is an even function of the variable $x$. In this section we will derive the general expressions without making further assumptions about the functional form of $f$ and defer a discussion of reasonable choices for $f$ to a later section.
@@
We have now obtained the results we need to coarse-grain the Madelung equations. In all that follows we assume an isotropic finite scale coarse-graining with the conventions defined above.
@@ III.2 Momentum Equation
The quantum stress is simply the Favre average of the quantum pressure term in the Madelung equations, the interaction stress is, similarly, the Favre average of the term due to interactions. The classical stress term has the form of the divergence of a stress tensor. In principle, one could perform the averages in Eq. (17) for specific quantum systems to study how these terms depend on the macroscopic flow variables. However, such a study is beyond the scope of the current work. Instead, following the finite scale analysis for classical fluids, we will expand each term in Eq. (17) perturbatively in $\ell^{2}$.
@@ III.3 Vorticity Equation
Note that, just as in classical hydrodynamics, the left hand side is the Lagrangian derivative of the vorticity field, that is the time derivative of the vorticity in the frame of the fluid. Also, just as in classical hydrodynamics, a vortex-stretching term, $\boldsymbol{\omega}\cdot\boldsymbol{\nabla}\ \langle\langle\textbf{u}\rangle\rangle$, emerges on the right hand side, implying that vorticity will, generally, be enhanced when the velocity field increases along the direction parallel to $\boldsymbol{\omega}$. This phenomenon is thought to play an important role in the cascade of energy from larger to smaller length scales in classical turbulence. Since it arises naturally upon coarse-graining, it stands to reason that similar phenomena should be expected to play a role in macroscopic effective descriptions of quantum turbulence as well, even though vorticity is, strictly-speaking, absent in quantum fluids.
@@ IV Example Flow Field: A Line Vortex
To better understand the emergence of vorticity in the finite scale theory outlined in the previous section, we will now consider an example flow field which is chosen to be amenable to a semi-analytic treatment while also exhibiting emergent vorticity when coarse-grained.
With these motivations in mind, we choose the velocity field to be that associated with a classical line vortex and a density given by a Gaussian distribution in three dimensions. The wavefunction associated with these choices is given by:
$\psi_{0}=\frac{1}{(2\pi)^{3/4}\sigma^{3/2}}e^{-\frac{\textbf{x}\cdot\textbf{x}}{4\sigma^{2}}}e^{i\frac{m}{\hbar}\frac{\Gamma}{2\pi}\tan^{-1}\left(\frac{x_{2}}{x_{1}}\right)}.$ (27)
We can see that $\psi$ is normalized in space, since we have chosen $\psi_{0}^{*}\psi_{0}$ to be a normalized Gaussian distribution. While it is not a solution of the Schrödinger equation, it is a valid initial wavefunction which would evolve in time according to the Schrödinger equation and it is a relatively simple example of a wavefunction whose coarse-grained hydrodynamic description can be shown to possess finite vorticity.
@@
$\rho$ $=\frac{m}{(2\pi)^{3/2}\sigma^{3}}e^{-\frac{\textbf{x}\cdot\textbf{x}}{2\sigma^{2}}},$ (28)
u $=\frac{\Gamma}{2\pi}\frac{1}{x_{1}^{2}+x_{2}^{2}}\left[-x_{2}\hat{x}_{1}+x_{1}\hat{x}_{2}\right],$
where $\Gamma$ is the circulation around the line vortex. We note that, while this flow field does possess finite circulation, given by $\Gamma$, it is actually irrotational in that $\boldsymbol{\nabla}\times\textbf{u}=0$ everywhere, except at the singular point $x_{1}=x_{2}=0$, as one can easily verify from the expression above.
@@
$\left\langle\rho\right\rangle$ $=\frac{m\prod_{i=1}^{3}\int_{-\infty}^{\infty}dx^{\prime}_{i}\exp\left[-\frac{\sigma^{2}x_{i}^{\prime 2}+\ell^{2}\left(x_{i}^{\prime}+x_{i}\right)^{2}}{2\sigma^{2}\ell^{2}}\right]}{(2\pi)^{3}\sigma^{3}\ell^{3}},$
$=\frac{m\exp\left\{-\frac{\textbf{x}\cdot\textbf{x}}{2\left(\sigma^{2}+\ell^{2}\right)}\right\}}{(2\pi)^{3}}\left(\frac{2}{\sigma^{2}+\ell^{2}}\right)^{3/2}$
$\times\prod_{i=1}^{3}\int_{-\infty}^{\infty}du_{i}\exp\left[-\left(u_{i}+\frac{\ell x_{i}}{\sigma\sqrt{2\left(\sigma^{2}+\ell^{2}\right)}}\right)^{2}\right],$
performing the final Gaussian integrals we have:
@@
$\left\langle\rho\right\rangle=\frac{m\exp\left\{-\frac{\textbf{x}\cdot\textbf{x}}{2\left(\sigma^{2}+\ell^{2}\right)}\right\}}{\left[2\pi\left(\sigma^{2}+\ell^{2}\right)\right]^{3/2}}.$ (30)
As expected, in the limit of $\ell\rightarrow 0$ this equals $\rho$.
This allows us to write the velocity field as:
@@
$\left\langle\left\langle u_{1}\right\rangle\right\rangle$ $=-\frac{\Gamma\sqrt{\ell^{-2}+\sigma^{-2}}}{2^{3/2}\pi^{2}}\int_{-\infty}^{\infty}d\zeta_{1}\int_{-\infty}^{\infty}d\zeta_{2}$ (31)
$\times\zeta_{2}\frac{\exp\left[-\left(\zeta_{1}-\frac{x_{1}}{\alpha}\right)^{2}-\left(\zeta_{2}-\frac{x_{2}}{\alpha}\right)^{2}\right]}{\zeta_{1}^{2}+\zeta_{2}^{2}},$
@@
$\left\langle\left\langle u_{2}\right\rangle\right\rangle$ $=\frac{\Gamma\sqrt{\ell^{-2}+\sigma^{-2}}}{2^{3/2}\pi^{2}}\int_{-\infty}^{\infty}d\zeta_{1}\int_{-\infty}^{\infty}d\zeta_{2}$
$\times\zeta_{1}\frac{\exp\left[-\left(\zeta_{1}-\frac{x_{1}}{\alpha}\right)^{2}-\left(\zeta_{2}-\frac{x_{2}}{\alpha}\right)^{2}\right]}{\zeta_{1}^{2}+\zeta_{2}^{2}},$
@@
$\left\langle\left\langle u_{3}\right\rangle\right\rangle$ $=0,$
where we define the constant $\alpha\equiv\ell\sqrt{2\left(\sigma^{2}+\ell^{2}\right)}/\sigma$. Since the purpose of this example is to demonstrate the presence of finite vorticity, we will compute $\boldsymbol{\omega}$ directly without explicitly evaluating $\langle\langle\textbf{u}\rangle\rangle$. This can be done by taking the curl of the velocity defined above:
$\boldsymbol{\omega}$ $=\boldsymbol{\nabla}\times\langle\langle\textbf{u}\rangle\rangle=\hat{x}_{3}\left[\partial_{x_{1}}\left\langle\left\langle u_{2}\right\rangle\right\rangle-\partial_{x_{2}}\left\langle\left\langle u_{1}\right\rangle\right\rangle\right]=\hat{x}_{3}\omega,$
where $\omega$ is given by the integral:
$\omega(x_{1},x_{2})$ $=\frac{\Gamma}{2\pi^{2}\ell^{2}}\left\{\pi-\int_{-\infty}^{\infty}d\zeta_{1}\int_{-\infty}^{\infty}d\zeta_{2}\right$. (32)
$\left.\times\frac{\zeta_{1}\frac{x_{1}}{\alpha}+\zeta_{2}\frac{x_{2}}{\alpha}}{\zeta_{1}^{2}+\zeta_{2}^{2}}e^{-\left(\zeta_{1}-\tfrac{x_{1}}{\alpha}\right)^{2}-\left(\zeta_{2}-\tfrac{x_{2}}{\alpha}\right)^{2}}\right\}.$
Defining the radius from the vortex core as $r=\sqrt{x_{1}^{2}+x_{2}^{2}}$ we can write the vorticity just as a function of $r$:
$\omega(r)$ $=\frac{\Gamma}{2\pi^{2}\ell^{2}}\left\{\pi-\frac{r}{\alpha}\int_{0}^{\infty}dr^{\prime}\int_{0}^{2\pi}d\phi\right$. (33)
$\left.\times e^{-r^{\prime 2}+\frac{2r}{\alpha}r^{\prime}\cos\phi}e^{-\frac{r^{2}}{\alpha^{2}}}\cos\phi\right\},$
$=\frac{\Gamma}{2\pi\ell^{2}}\left\{1-\frac{1}{2\sqrt{\pi}}\frac{r}{\alpha}\int_{0}^{2\pi}d\phi\right$.
$\left.\times e^{-\frac{r^{2}}{\alpha^{2}}\sin^{2}\phi}\cos\phi\ \text{erf}\left(\frac{r\cos\phi}{\alpha}\right)\right\}.$
Written in this form, it is clear that, along the center line of the vortex $r=0$, the vorticity is equal to the vortex strength divided by the area $\ell^{2}$:
$\omega_{0}=\frac{\Gamma}{2\pi\ell^{2}}.$ (34)
Figure 1: Plot of emergent vorticity as given by Eq. (33) as a function of radius from vortex center. Vorticity is normalized to $\omega_{0}$ as defined in Eq. (34) and radii are expressed in units of $\alpha=\ell\sqrt{2\left(\sigma^{2}+\ell^{2}\right)}/\sigma$.
Further insight can be gained by expanding the error function using its Taylor series representation:
$\omega(r)$ $=\omega_{0}\left\{1-\frac{1}{\pi}\sum_{n=0}^{\infty}\left(\frac{r}{\alpha}\right)^{2(n+1)}\int_{0}^{2\pi}d\phi\right$.
$\left.\times e^{-\frac{r^{2}}{\alpha^{2}}\sin^{2}\phi}\left(\cos\phi\right)^{2(n+1)}\frac{\left(-1\right)^{n}}{n!\left(2n+1\right)}\right\}.$
Taking the leading order term in $r/\alpha$ we find that for $r<<\alpha=\ell\sqrt{2\left(\sigma^{2}+\ell^{2}\right)}/\sigma$ the vorticity is given approximately by:

### v2

@@ Abstract
The Madelung equations offer a hydrodynamic description of quantum systems, from single particles to quantum fluids. In this formulation, the probability density is mapped onto the fluid density and the phase is treated as a scalar potential generating the velocity field. As examples of potential flows, quantum fluids described in this way are inherently irrotational, but quantum vortices may arise at discrete points where the phase is undefined. In this paper, starting from this irrotational description of a quantum fluid, a coarse-graining procedure is applied to arrive at a macroscopic description of the quantum fluid in terms of a hierarchy of moments in which the role of velocity is played by a Favre average of the microscopic velocity field. This hierarchy is truncated using an explicit closure derived from an expansion in a finite length scale. The resulting coarse-grained fields are shown to allow for finite vorticity at any point in the fluid. Furthermore, it is shown that this vorticity obeys a similar equation to the vorticity equation in classical hydrodynamics and includes a vortex-stretching term. The particular closure employed here also gives rise to a novel stress term in the fluid equations, which in the appropriate limit appears analogous to an artificial viscous stress from computational fluid dynamics.
@@
Fluid turbulence is known to be a ubiquitous non-equilibrium phenomenon arising in fluids when the kinetic energy injected at large scales dominates the dissipation of energy at small scales due to viscosityMcComb (1990, 1995); Sreenivasan (1999); Eyink and Sreenivasan (2006). While first identified and extensively studied in classical fluids, turbulence has also been observed in quantum fluids, including both superfluids and ultra-cold atomic gasesVinen and Niemela (2002); Barenghi and Donnelly (2009); Vinen (2010); Paoletti and Lathrop (2011); Tsubota (2014); Barenghi et al. (2014). While there is no consensus on the precise definition of turbulence, it is characterized by spatially complex and temporally aperiodic flow fields and involves processes on many length scales. In three-dimensions, it is well established that turbulence involves a cascade of energy from large to small length scalesKolmogorov (1991); Falkovich (2009); Dubrulle (2019); Carbone and Bragg (2020). While the details of this process are not fully understood, there is significant evidence that, in classical turbulence, this cascade is related to the non-linear phenomena of strain self-amplification and vortex stretchingCarbone and Bragg (2020). In quantum turbulence, this cascade has a remarkably similar behavior at large length scales, with the Kolmogorov spectrum being observed in superfluid 4He and simulations using the Non-linear Schrödinger equation (NLSE)Paoletti and Lathrop (2011). However, in contrast to the case of classical turbulence where the cascade is related to the dynamics of a continuous vorticity field, the cascade in quantum turbulence is related to the dynamics of a discrete number of quantized vortices. The differences between these two cases are most pronounced at length scales less than the inter-vortex spacing, where significant contributions to the energy cascade have their origin in transverse modes (Kelvin waves) which propagate along the quantized vorticesBarenghi and Donnelly (2009). The two cases appear much more similar at larger length scales where the dynamics are dominated by the collective motion of many vortices. While this similarity between classical and quantum turbulence is not unexpected based on intuition from the correspondence principlePaoletti and Lathrop (2011), a full theory establishing the rigorous connection between these two regimes is lacking, due in part to the lack of a single governing equation used to describe quantum turbulence at all length scales.
@@
In contrast to the study of classical turbulence, whose phenomenology is believed to emerge from the dynamics governed by a single equation, the Navier-Stokes equation, current descriptions of quantum turbulence require a hierarchy of modelsBarenghi et al. (2014). The microscopic behavior of the quantum fluid is modeled using the NLSE, or appropriate generalizations to capture strong interactions; the mesoscale behavior is described by vortex-filament models; and macroscopic dynamics are described using the Hall–Vinen–Bekharevich–Khalatnikov (HVBK) modelBarenghi et al. (2014); Zhang et al. (2023). Previous works have investigated the coarse-grained dynamics of many-vortex solutions of the GPE, either by deriving an emergent non-Eulerian vortex-fluid hydrodynamics from the point-vortex modelYu and Bradley (2017), or by developing kinetic and statistical theories based on coarse-grained vortex distributionsSalman and Maestrini (2016). The quantum fluid equations have also been coarse-grained directly, in a work by Tanogami, using a spatial filter to derive a hierarchy of filtered moment equationsTanogami (2021). A key purpose of the present paper is to extend the coarse-graining framework introduced by Tanogami by providing a closure of the hierarchy of filtered moment equations. This closure yields explicit expressions for the emergent sub-scale stress terms that appear in the unclosed formulation. In addition, we demonstrate, using a family of simple analytic examples, that the coarse-grained velocity field in three dimensions can exhibit finite, continuous vorticity, and that this coarse-grained vorticity obeys an evolution equation closely analogous to the classical vorticity equation. The emergence of a novel viscous stress term due to integrating out small scale degrees of freedom has similarities with a recent work which demonstrated a relationship between decoherence and viscosityZhou et al. (2025).
@@
The particular closure introduced in this paper is based on finite scale theory, a method previously studied in the context of classical continuum dynamicsMargolin and Rider (2002); Margolin (2009, 2014, 2019). The motivation for finite scale theory in classical continua is based on the fact that in order to solve many problems in classical continuum dynamics, the continuum equations are almost always approximated by discretizing the domain and forming a mesh, represented by a finite number of points in the original space. In the context of Lagrangian hydrocodes, this discretization amounts to an averaging of the continuum variables over finite size cells. However, these averaged quantities are not necessarily solutions to the original partial differential equations governing the dynamics of the continuum. Finite scale theory provides a connection between the coarse-graining of the continuous domain and the differential equations describing the averaged quantities.
The remainder of this paper is organized as follows. In Sec. I the quantum fluid equations are introduced explicitly, highlighting their similarities to classical fluid equations. Sec. II contains a pedagogical derivation of the coarse-graining procedure which will be applied to the quantum fluid equations as well as a derivation of the closure from finite scale theory which will be used to truncate the hierarchy of filtered moments. In Sec. III the coarse-graining procedure and closure model are applied to the Madelung equations and both the general (unclosed) and finite-scale (closed) equations are presented and discussed. In Sec. III.3 the equation governing the evolution of vorticity is derived based on the results in the previous sections. Sec. IV presents a family of simple examples in which an irrotational velocity field gives rise to a finite and continuous vorticity field after undergoing the coarse-graining procedure discussed in the preceding sections. Sec. V provides a discussion of the conclusions that can be drawn from both the general equations as well as the simple examples.
@@ I Madelung Equations
$\frac{1}{\rho}\boldsymbol{\nabla}\cdot\boldsymbol{\sigma}+\textbf{f}=\frac{\hbar^{2}}{2m^{2}}\boldsymbol{\nabla}\left(\frac{\nabla^{2}\sqrt{\rho}}{\sqrt{\rho}}\right)-\frac{g}{m}\boldsymbol{\nabla}\rho.$
@@ II Generalized Finite Scale Analysis
where $f$ is a normalized distribution describing the details of our coarse-graining procedure. At this level, the spatial filtering procedure here is identical to the one employed by TanogamiTanogami (2021) in his analysis of the energy cascade in quantum fluids.
The procedure employed in previous studies of finite scale closures would correspond to the choice of a uniform box distributionMargolin and Rider (2002); Margolin (2009); however, in principle, many other choices could yield similar results. For simplicity, we will assume that $f$ is a separable distribution, $f(\textbf{x})=f_{1}(x_{1})f_{2}(x_{2})f_{3}(x_{3})$ and that each function $f_{i}(x)$ is an even function of the variable $x$. In this section we will derive the general expressions without making further assumptions about the functional form of $f$ and defer a discussion of reasonable choices for $f$ to a later section.
@@
As we will demonstrate in the next section, Eqs. (9) provide sufficient conditions to close the coarse-grained Madelung equations, representing an extension of the coarse-graining framework employed in previous workTanogami (2021).
@@ III.2 Momentum Equation
The quantum stress is simply the Favre average of the quantum pressure term in the Madelung equations, the interaction stress is, similarly, the Favre average of the term due to interactions. The classical stress term has the form of the divergence of a stress tensor. These terms are consistent with previous analyses of the Madelung equationsTanogami (2021). In principle, one could perform the averages in Eq. (17) for specific quantum systems to study how these terms depend on the macroscopic flow variables. However, such a study is beyond the scope of the current work. Instead, following the finite scale analysis for classical fluids, we will expand each term in Eq. (17) perturbatively in $\ell^{2}$, thus extending the unclosed hierarchy of moments previously studied to a closed set of equations for coarse-grained quantities.
@@ III.3 Vorticity Equation
Note that, just as in classical hydrodynamics, the left hand side is the Lagrangian derivative of the vorticity field, that is the time derivative of the vorticity in the frame of the fluid. Also, just as in classical hydrodynamics, a vortex-stretching term, $\boldsymbol{\omega}\cdot\boldsymbol{\nabla}\ \langle\langle\textbf{u}\rangle\rangle$, emerges on the right hand side, implying that vorticity will, generally, be enhanced when the velocity field increases along the direction parallel to $\boldsymbol{\omega}$. This phenomenon is thought to play an important role in the cascade of energy from larger to smaller length scales in classical turbulence. Since it arises naturally upon coarse-graining, it stands to reason that similar phenomena should be expected to play a role in macroscopic effective descriptions of quantum turbulence as well, even though vorticity is, strictly-speaking, absent in quantum fluids. The emergence of this phenomenon is consistent with Tanogami’s previous work studying the energy cascade in coarse-grained quantum fluidsTanogami (2021) and lends further credence to his hypothesis.
@@ IV Example Flow Field: A Line Vortex
To better understand the emergence of vorticity in the finite scale theory outlined in the previous section, we will now consider an example flow field which is chosen to be amenable to an exact analytic treatment while also exhibiting emergent vorticity when coarse-grained.
With these motivations in mind, we choose the velocity field to be that associated with a classical line vortex and a density away from the vortex core given by a Gaussian distribution in three dimensions. At a distance $R$ from the vortex core, we give the density a simple scaling $\rho\sim R^{2n}$, for $n\in\mathbb{N}$. The wavefunction associated with these choices is given by:
$\psi_{0}=\frac{\left(\textbf{x}_{\perp}\cdot\textbf{x}_{\perp}\right)^{n/2}}{(2\pi)^{3/4}\sigma^{3/2}r_{0}^{n}}e^{-\frac{\textbf{x}\cdot\textbf{x}}{4\sigma^{2}}}e^{i\frac{m}{\hbar}\frac{\Gamma}{2\pi}\tan^{-1}\left(\frac{x_{2}}{x_{1}}\right)},$ (27)
where $\textbf{x}_{\perp}$ lies in the plane perpendicular to the vortex line and $r_{0}$ is chosen to normalize the wavefunction:
$r_{0}=\sqrt{2}\sigma\left(n!\right)^{\frac{1}{2n}}.$ (28)
While this is not, in general, a solution of the Schrödinger equation it is a valid initial wavefunction which would evolve in time according to the Schrödinger equation and it is a relatively simple example of a wavefunction which captures some of the essential features of a quantized vortex, and whose coarse-grained hydrodynamic description can be shown to possess finite vorticity.
@@
$\rho$ $=\frac{m\left(\textbf{x}_{\perp}\cdot\textbf{x}_{\perp}\right)^{n}}{(2\pi)^{3/2}2^{n}\sigma^{2n+3}n!}e^{-\frac{\textbf{x}\cdot\textbf{x}}{2\sigma^{2}}},$ (29)
u $=\frac{\Gamma}{2\pi}\frac{-x_{2}\hat{x}_{1}+x_{1}\hat{x}_{2}}{\textbf{x}_{\perp}\cdot\textbf{x}_{\perp}},$
where $\Gamma$ is the circulation around the line vortex. We note that, while this flow field does possess finite circulation, given by $\Gamma$, it is actually irrotational in that $\boldsymbol{\nabla}\times\textbf{u}=0$ everywhere, except at the singular point $x_{1}=x_{2}=0$, as one can easily verify from the expression aboveBatchelor (2000).
@@
We can now evaluate the coarse-grained fields, $\left\langle\rho\right\rangle$ and $\langle\langle\textbf{u}\rangle\rangle$. The details of this calculation, for generic integer $n\geq 0$, are given in Appendix A. For the coarse-grained density we find, Eq. (47):
@@
$\left\langle\rho\right\rangle=e^{-\frac{z^{2}}{2\ell^{2}}\left(1-\frac{\tilde{\sigma}^{2}}{\ell^{2}}\right)}\frac{e^{-\frac{R^{2}}{2\ell^{2}}}m\left(2\tilde{\sigma}^{2}\right)^{n+3/2}}{(4\pi)^{3/2}\ell^{3}2^{n}\sigma^{2n+3}}M\left(n+1,1,\kappa^{2}\right),$ (31)
and for the coarse-grained velocity we find, Eq. (48):
@@
$\left\langle\left\langle\textbf{u}\right\rangle\right\rangle=\frac{R\ \Gamma}{4\pi\ell^{2}}\frac{M\left(n+1,2,\kappa^{2}\right)}{M\left(n+1,1,\kappa^{2}\right)}\ \hat{\theta}$, (32)
where $M(a,b,z)$ is Kummer’s confluent hypergeometric function, $\tilde{\sigma}=\sigma\ell/\sqrt{\ell^{2}+\sigma^{2}}$, $\kappa^{2}=R^{2}\tilde{\sigma}^{2}/2\ell^{4}$, and $\hat{\theta}$ is the unit vector in the azimuthal direction around the vortex core.
From Eq. (32) it is straightforward to arrive at the vorticity, Eq. (54):
$\boldsymbol{\omega}=\hat{z}\frac{\Gamma}{2\pi\ell^{2}}\left\{1-\kappa^{2}\frac{M\left(n+1,2,\kappa^{2}\right)}{M\left(n+1,1,\kappa^{2}\right)}\left[1+n\frac{M\left(n+1,2,\kappa^{2}\right)}{M\left(n+1,1,\kappa^{2}\right)}\right]\right\},$ (33)
where $\hat{z}$ is in the direction along the vortex line. This demonstrates the emergence of vorticity under the coarse-graining procedure described in the preceding sections, and Eq. (33) can be used to numerically compute the vorticity for arbitrary $n\in\mathbb{N}$. However, much insight can be gained from considering the special cases of $n=0$ and $1$.
As we show in Appendix A.1, the coarse-grained variables for $n=0$ can be written as:
$\left\langle\rho\right\rangle=e^{-\frac{z^{2}+R^{2}}{2\ell^{2}}\left(1-\frac{\tilde{\sigma}^{2}}{\ell^{2}}\right)}\frac{m\left(2\tilde{\sigma}^{2}\right)^{3/2}}{(4\pi)^{3/2}\ell^{3}\sigma^{3}},$ (34)
@@
$\left\langle\left\langle\textbf{u}\right\rangle\right\rangle=\frac{\Gamma}{2\pi}\frac{1-e^{-\frac{R^{2}\tilde{\sigma}^{2}}{2\ell^{4}}}}{R\sigma^{2}}\left(\sigma^{2}+\ell^{2}\right)\ \hat{\theta}$, (35)
and
$\boldsymbol{\omega}=\hat{z}\frac{\Gamma}{2\pi\ell^{2}}e^{-\frac{R^{2}\tilde{\sigma}^{2}}{2\ell^{4}}}.$ (36)
From Eq. (36) we see that the magnitude of the vorticity at the vortex core, $R=0$, is $\omega=\Gamma/2\pi\ell^{2}$. In the limit of $\ell\rightarrow 0$ this diverges, consistent with the singular nature of the microscopic fields. Away from the vortex core, $R>0$, we see that $\lim_{\ell\rightarrow 0}\omega=0$, consistent with the irrotational nature of the microscopic velocity field. However, we note that for $n=0$ the density does not vanish at the vortex core, and so this case is not as physically relevant as the $n=1$ case.
As we show in Appendix A.2, the coarse-grained variables for $n=1$ can be written as:
$\left\langle\rho\right\rangle=e^{-\frac{z^{2}+R^{2}}{2\ell^{2}}\left(1-\frac{\tilde{\sigma}^{2}}{\ell^{2}}\right)}\frac{m\left(2\tilde{\sigma}^{2}\right)^{5/2}}{(4\pi)^{3/2}\ell^{3}2\sigma^{5}}\left(1+\frac{R^{2}\sigma^{2}}{2\ell^{2}\left(\ell^{2}+\sigma^{2}\right)}\right),$ (37)
@@
$\left\langle\left\langle\textbf{u}\right\rangle\right\rangle=\frac{R\ \Gamma}{2\pi}\frac{1}{2\ell^{2}+\tfrac{R^{2}\sigma^{2}}{\ell^{2}+\sigma^{2}}}\ \hat{\theta}$, (38)
and